[Hallc_running] [New Logentry] RC: Daily Update

mdiefent at jlab.org mdiefent at jlab.org
Tue Dec 18 10:15:02 EST 2018

Logentry Text:
<p><strong><span style="font-family: Helvetica; font-size: 12px;">Update on day and swing shifts of Monday</span><span style="font-family: Helvetica; font-size: 12px;">, December 17, and owl shift of Tuesday, December 18</span></strong><br style="font-family: Helvetica; font-size: 12px;"><br style="font-family: Helvetica; font-size: 12px;"><strong><span style="font-family: Helvetica; font-size: 12px;">CEBAF status</span></strong><span style="font-family: Helvetica; font-size: 12px;">:</span><span style="font-family: Helvetica; font-size: 12px;"> </span></p><div style="font-family: Helvetica; font-size: 12px;"><ul><li>Beam studies from 8:00 - 11:00 a.m.. Received in the remaining day shift 4.5 hours of beam (https://logbooks.jlab.org/entry/3639765). Jay Benesch has studied the effect of the SHMS fringe fields on the beam (https://logbooks.jlab.org/entry/3639736). </li><li>Received 7.8 hours of beam during the swing shift (https://logbooks.jlab.org/entry/3639875). </li><
 li>Received 7.6 hours of beam during the owl shift (https://logbooks.jlab.org/entry/3639964). </li><li>The physics run will end on Thursday, December 20, at 6:30 a.m. to prepare the accelerator shutdown at 7:00 a.m.. The accelerator and experimental facilities will be put into restricted access. </li></ul></div><p><strong><span style="font-family: Helvetica; font-size: 12px;">Hall C status</span></strong><span style="font-family: Helvetica; font-size: 12px;">: </span></p><div style="font-family: Helvetica; font-size: 12px;"><ul><li>Three-pass beam with an energy of 4.933 GeV is delivered to Hall C. </li><li>Hall C is setup for the Kaon LT experiment (E12-09-011). The experiments aims for the first precision measurements of L/T separated kaon electroproduction cross sections above the resonance region. Details on the Kaon LT experiment can be found on: https://redmine.jlab.org/projects/kltexp/wiki</li><li>During the end of the owl shift on December 16, the co
 llimator position in the "HMS Collimator Mot!
 ion Control" GUI was shown as "home" instead of "large collimator" (https://logbooks.jlab.org/entry/3639324). The displayed position did not update after the reset and homing procedure (https://logbooks.jlab.org/entry/3639344). Dave Gaskell has compared xptar/yptar, "xsieve", "ysieve" and other variables between runs recorded in the morning of December 16 and runs recorded on December 15 and verified that the collimator position has not changed (https://logbooks.jlab.org/entry/3639351). So we decided to continue data taking. During the controlled access, Ethan Becker inspected the HMS collimator motion control. As any work on the control and its readout would require a move of the collimator, we decided to postpone the work after the end of data taking with a beam energy of 4.933 GeV (https://logbooks.jlab.org/entry/3639672).</li><li>On Monday, December 17, we changed in a controlled access the SHMS angle to 6.00 degree (https://logbooks.jlab.org/entry/3639656).&
 nbsp;</li></ul></div><p><strong><span style="font-family: Helvetica; font-size: 12px;">Data taking</span></strong><span style="font-family: Helvetica; font-size: 12px;">: </span></p><div style="font-family: Helvetica; font-size: 12px;"><ul><li>The detailed run plan is available on: https://redmine.jlab.org/attachments/download/590/runplan_4p9gev.pdf</li><li>We have done the harp scan: https://logbooks.jlab.org/entry/3638031 (see also 3638030 and the work documented in entries 3638025 and 638027)</li><li>We have completed the energy measurement during the day shift of December 14: https://logbooks.jlab.org/entry/3638789 (and 3638787 for details on beam size). The beam energy is 4.9320 +/- 0.0026 GeV. </li><li>We have taken the data for the calibration runs: Holly has done the sieve and beam position comparisons: https://logbooks.jlab.org/entry/3638085. Richard documented his analysis of the heep coincidence runs: https://log
 books.jlab.org/entry/3638592 Nooffsets !
 are obser!
 ved in the missing mass and momentum distributions. </li><li>We are taking data for the SHMS angle of 6.00 degree with a beam charge of 16uA and LH2 target. On Tuesday, December 18, we will collect data for the thick dummy target. On Wednesday, December 19, we will decide if we will take data also for smaller beam charges than 16uA. We have completed our charge goals for the SHMS angles of 8.86 and 11.86 degrees and for both LH2 and dummy targets. </li></ul></div><p><strong><span style="font-family: Helvetica; font-size: 12px;">Hall C Daily RC meeting</span></strong><span style="font-family: Helvetica; font-size: 12px;">: The next meeting will be on Wednesday, December 19.</span><span style="font-family: Helvetica; font-size: 12px;"> </span></p>


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3640020
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181218/e981d6d3/attachment.html>

More information about the Hallc_running mailing list